# How General Are Measures of Choice Consistency? Evidence from Experimental and Scanner Data

### v1

@@ Abstract
Choice consistency with utility maximization is a fundamental assumption in economic analysis and is extensively measured across various contexts. Here we investigate the generalizability of consistency measures derived from purchasing decisions using supermarket scanner data and budgetary decisions from lab-in-the-field experiments. We observe a lack of correlation between consistency scores from supermarket purchasing decisions and those from risky decisions in the experiment. However, we observe moderate correlations among experimental tasks and low to moderate correlations across purchasing categories and time periods within the supermarket. These results suggest that choice consistency may be characterized as a multidimensional skill set.
Keyword: consumer choice, revealed preference, choice consistency, utility maximization, lab-in-the-field experiment JEL: C91, D81, D91
@@ 1 Introduction
Central to economic analysis is the assumption that a decision maker (DM) maximizes her utility function given her budget constraint. Revealed preference analysis characterizes the conditions under which a DM, based on a given choice dataset, indeed maximizes a well-behaved utility function (Samuelson, 1938; Afriat, 1967; Varian, 1982). It thus provides a powerful framework and serves as a textbook example of normative approaches on how decisions should be made. Going beyond normative approach, researchers have examined the descriptive validity of this framework, and used it to assess choice consistency in various settings (Crawford and De Rock, 2014; Chambers and Echenique, 2016). A growing body of literature has measured consistency scores in risky, intertemporal, and social budgetary decisions in experiments, as well as in purchasing decisions derived from expenditure surveys and from scanner data in supermarkets. [^1] For studies utilizing experimental data: Ahn et al. (2014); Andreoni and Miller (2002); Choi et al. (2007); Echenique et al. (2023); Fisman et al. (2007); Halevy et al. (2018); Polisson et al. (2020); Dembo et al. (2024); Ellis and Freeman (2024), and for studies employing purchasing data: Blundell et al. (2003, 2008); Crawford (2010); Cherchye et al. (2020); Dean and Martin (2016); Echenique et al. (2011). Consistency with utility maximization is often referred to as choice consistency or rationality. For simplicity, we use choice consistency throughout this paper. Furthermore, it has been suggested that consistency scores reflect an individual’s capacity to make sound decisions and may serve as a significant determinant of wealth disparities among individuals in the general population (Choi et al., 2014) and of development gaps between countries (Cappelen et al., 2023).
@@
Building on these insights into the notion of choice consistency and its applicability across settings, here we take a closer look at its empirical validity and test its generalizability—the extent to which DMs exhibiting high consistency scores in one setting also demonstrate high scores in another setting. More specifically, do DMs with high consistency scores as consumers in a supermarket environment also have higher scores as participants in a controlled experimental setting? Two related questions arise: Do DMs show correlated consistency scores across risk, social, and food-related budgetary decisions in the experiment? Do DMs exhibit correlated consistency scores across different consumption categories and time periods in the supermarket? On the one hand, it is plausible that measures of choice consistency can be generalized across settings, as they may reflect a DM’s overall capacity of making good decisions. Conversely, it is also conceivable that these measures are setting-specific, as DM may possess different preference structures, face varying budget constraints, and have distinct experiences or decision rules across settings, leading to divergent scores inferred from choice data.
@@
We address these questions by combining experimental and scanner data. First, we examine consistency scores of the same individuals who make risky decisions in the lab-in-the-field experiment and make consumption decisions in the supermarket. Using the scanner data, we restrict our attention to 6,144 consumers of whom we have purchase records of meat and vegetables—the two most common consumption categories—over 24 consecutive months, so that we have sufficient power for the revealed preference analysis (see Section 3). We then successfully invite 1,055 of these 6,144 consumers to participate in a budgetary task (Experiment 1) in which participants make 22 portfolio decisions between two Arrow securities with different budget lines (Choi et al., 2007, 2014; Halevy et al., 2018; Kim et al., 2018). They need to allocate 100 experimental tokens between two accounts in which tokens are converted to cash with different exchange rates, and the amount in one of the two accounts will be paid out randomly with 50 percent probability. Therefore, by extracting these 1,055 consumers from the scanner data and aggregating the consumption at the month level, we obtain the choice dataset in the supermarket (Scanner Dataset 1).
@@
We measure the choice consistency by evaluating the extent to which choices adhere to the Generalized Axiom of Revealed Preference (GARP), a necessary and sufficient condition whereby a dataset can be rationalized by a well-behaved utility function (in accordance with utility maximization). To assess how closely an individual’s choice dataset complies with GARP, we compute the commonly used critical cost efficiency index (CCEI, Afriat, 1972), as well as the Houtman-Maks index (HMI, Houtman and Maks, 1985); money pump index (MPI, Echenique et al., 2011); and minimum cost index (MCI, Dean and Martin, 2016). Based on Experiment 1 and Scanner Dataset 1, our primary finding is that consistency scores are uncorrelated between risky decisions in the experiment and food consumption decisions in the supermarket regardless of the index used.
@@
Second, we test whether the observed near-zero correlation is attributed to the differences in the types of decisions—risky decisions versus food consumption decisions—or due to the disparities between experimental and supermarket settings, and conduct a second experiment (Experiment 2). We invite another 302 consumers to participate in three budgetary tasks, including the risk task used in Experiment 1, social task (Andreoni and Miller, 2002; Fisman et al., 2007) and food task (Chen et al., 2023; Harbaugh et al., 2001). We find that consistency scores are moderately correlated across these three tasks, where the magnitude of correlations is the highest between risk and social tasks.
@@
Third, we also examine the correlation of consistency scores for choices within supermarket using Scanner Dataset 2. The dataset consists of 822 consumers across different consumption categories including meat, vegetables, fruits and snacks, and 938 consumers across different time periods from 2018 to 2021. We find that the correlations of consistency scores are generally low for choices across consumption categories and over different time periods within supermarket.
@@
We perform several robustness checks for these observed correlations. We use alternative indices of HMI, MPI, and MCI, control the power of detecting GARP violations, impose additional assumptions on preference structures using revealed preference analysis, and assess the fit of structural estimations of preferences. Our findings remain consistent across these variations with two exceptions. Namely, when we impose more assumptions on preference structures or evaluate the fit in structural estimations, the correlations within scanner data improve, both across categories and over time periods.
@@
We investigate several factors that may influence consistency scores across different environments, including randomness in choice behaviors, budget constraints, formation of preferences, and the application of heuristic rules. Notably, our findings indicate that choices made in supermarkets tend to be more random compared to those made in experimental settings. Additionally, consistency in supermarket choices is influenced by various contextual factors, such as seasonality, time of day, promotional discounts, and individual shopping experiences. In contrast, consistency in experimental choices is shaped by learning effects across rounds of decisions, the use of heuristic rules, and participants’ cognitive abilities. These distinct elements play crucial roles in how individuals navigate decision-making in the supermarket and in the experiment, and highlight the multifaceted nature of decision-making across settings.
@@
Our study adds to the literature on measuring consistency with utility maximization based on revealed preference analysis. In addition to theoretical and experimental studies, the measures of choice consistency have been proposed as proxies for decision-making quality and widely used in applied settings, which link them to education, occupation, borrowing behavior of high-cost loan, income, and wealth (e.g., Banks et al., 2019; Cappelen et al., 2023; Carvalho et al., 2016, 2024; Choi et al., 2014; Kim et al., 2018; Li et al., 2023). In this literature, most studies focus on a specific type of decisions either in the experiment or in the field. One exception is Kim et al. (2018), in which they collect choice data on both risky and intertemporal decisions in the laboratory experiment to evaluate effects of an education program. Based on consistency scores of risky and intertemporal decisions in their dataset, we compute the Spearman’s rank correlation to be 0.51, which is in line with the observed correlations across tasks in our Experiment 2. Our study is the first to examine the generalizability of consistency measures, and contributes to a better understanding of these measures.
@@
Our study also contributes to the literature on the generalizability of choice behaviors across settings (Camerer, 2015; Levitt and List, 2007; Snowberg and Yariv, 2021; Chapman et al., 2024). A question that has attracted great attention is the generalizability of risk preferences. For example, Weber et al. (2002) show that risk attitudes vary across domains; these include financial, health/safety, recreational, ethical, and social decisions. Dohmen et al. (2011) find that different measures of risk attitudes are imperfectly correlated across settings. Barseghyan et al. (2011) use three deductible choices made in the domain of auto and homeowner insurance to separately estimate individual risk attitude using structural approach, and reject the null of fully domain-general risk attitudes. Einav et al. (2012) provide evidence that in support of positive correlations of risk preferences within five employer-provider insurance coverage decisions, and weaker relationships between these insurance decisions and investment decisions with respect to a 401(k) plan. In addition to risk preferences, moderate or weak correlations across settings have also been reported in time preferences (Augenblick et al., 2015; Burks et al., 2012); social preferences (Bruhin et al., 2019; Fehr and Leibbrandt, 2011; Galizzi and Navarro-Martinez, 2019); and strategic sophistication (Georganas et al., 2015; Levitt et al., 2011; Rubinstein, 2016). These studies have made important contributions to the extent to which theory-guided experimental findings can be generalized and applied to real-world scenarios, which is essential for theoretical development, experimental design, and practical application.
@@
In this study, we examine choice consistency using risky, social and food decisions in two experiments and consumption purchases in the supermarket. On the one hand, the lack of correlation between experimental and scanner data suggests that consistency scores revealed from choice behaviors in various settings may not solely reflect the general ability to make good decisions according to one’s preferences. On the other hand, we do observe moderate correlations across three tasks in the experiment, and some weak to moderate correlations across categories and time in the supermarket, which indicate that consistency measures can be generalized within settings.
@@
Apart from measurement and identification issues, these findings point to the possibility that choice consistency is multidimensional. This perspective resonates with developments in the literature on intelligence, which have evolved over decades from a one-dimensional understanding of intelligence (Spearman, 1904) to the recognition of two dimensions—fluid and crystallized intelligence (Cattell, 1943), and ultimately to the widely accepted view of multidimensional intelligence (McGrew, 2009; Horn and Cattell, 1966). In the economics literature, accumulative evidence has provided support to the notion that choice behaviors are interconnected and can be reduced to several underlying common factors (Chapman et al., 2023; Dean and Ortoleva, 2019; Stango and Zinman, 2023), and the ability to make good decisions is multidimensional (Deming, 2021). Building on these studies, our study takes an initial step to address an important yet unexplored question on the generalizability of choice consistency, and contributes to a deeper understanding of the complexities inherent in decision-making processes.
@@
The rest of the paper is organized as follows. Section 2 presents the theoretical background of revealed preference analysis. Section 3 describes the scanner data and experimental design. Section 4 presents our main results, and Section 5 examines the underlying mechanisms. Section 6 provides some concluding remarks.
@@ 2 Theoretical Background
Consider a DM who chooses bundles $x^{t}\in\mathbb{R}_{+}^{K}$ from budget lines {$x:p^{t}\cdot x\leq p^{t}\cdot x^{t},p^{t}\in\mathbb{R}_{++}^{K}$}. A dataset $\mathcal{O}=\left\{\left(p^{t},x^{t}\right)\right\}_{t=1}^{T}$ refers to a collection of $T$ decisions of the DM. Let $\mathcal{X}=\{x^{t}\}_{t=1}^{T}$ be the set of bundles chosen by the DM. We say that $x^{i}$ is directly revealed preferred to $x^{j},$ denoted by $x^{i}\succsim^{\ast}x^{j},$ if a DM chooses $x^{i}$ when $x^{j}\in\mathcal{X}$ is affordable (i.e., $p^{i}\cdot x^{j}\leq p^{i}\cdot x^{i}$). Denote the asymmetric part as $\succ^{\ast},$ which refers to the relation of directly strictly revealed preference. Denote $\succsim^{\ast\ast}$ the transitive closure of $\succsim^{\ast},$ which refers to the revealed preferred relation. A dataset $\mathcal{O}$ satisfies the Generalized Axiom of Revealed Preference (GARP) if the following holds:
@@ 2.1 Indices of Choice Consistency
Critical Cost Efficiency Index (CCEI)  A popular approach for measuring the departure from utility maximization is the critical cost efficiency index (CCEI) proposed by Afriat (1972). A DM has a CCEI $e\in\left[0,1\right]$ if $e$ is the largest number with a well-behaved $U$ that rationalizes the dataset for every $x^{t}\in\mathcal{X}:$
@@
A CCEI of 1 indicates passing GARP perfectly. A CCEI less than 1—say, 0.95—indicates that there is a utility function for which the chosen bundle $x^{t}$ is preferred to any bundle that is more than 5 percent cheaper than $x^{t}$. Put differently, the CCEI can be viewed as the amount by which a budget constraint must be relaxed in order to remove all violations of GARP, because the DM is spending more money than is required to achieve the utility targets (Polisson and Quah, 2024). Alternative, it may be linked to the behavioral notion of the just-noticeable difference, DM’s innate inability to distinguish similar bundles (Dziewulski, 2020).
@@
Houtman-Maks Index (HMI)  Houtman and Maks (1985) propose an alternative approach by measuring the maximal number of observations in the observed sample that are consistent with rational choices. For example, an HMI score of 0.941 indicates that the largest proportion of subset of choices consistent with GARP is 94.1 percent.
@@
Money Pump Index (MPI)  Echenique et al. (2011) provide an index to measure the amount of money one can extract from an individual for each violation of GARP. Their index is based on the idea that an individual with a GARP violation can be exploited by an “arbitrager” as a “money pump”. The “arbitrager” can choose the allocation $x^{1}$ at price $p^{2}$ and allocation $x^{2}$ at price $p^{1}$, then trade $x^{1}$ with the individual at $p^{1}$ and $x^{2}$ at $p^{2}$, which yields a profit of $p^{1}(x^{1}-x^{2})+p^{2}(x^{2}-x^{1})$. Given multiple violations of GARP, a money pump cost will be associated with each violation. Following Echenique et al. (2011), we use the mean money pump costs for cyclic sequences of allocations with the length of two. For example, an MPI score equal to 0.059 means on average 5.9 percent of expenditure can be exploited by an “arbitrager” from GARP violation.
Minimum Cost Index (MCI)  Dean and Martin (2016) propose MCI to measure the minimum cost of breaking all revealed preference cycles in a dataset. MCI is defined as $\min\limits_{B\subset R_{0}}\nicefrac{{\sum\limits_{(i,j)\in B}p^{i}(q^{i}-q^{j})}}{{\sum\limits_{t=1}^{T}p^{t}q^{t}}}$ such that $R_{0}/B$ is acyclic, where $R_{0}$ represents the set of preference relation $\succsim^{\ast\ast}$. The index has a high value when there are a large number of cycles for which all GARP violations are based on significant monetary differences relative to total expenditure. Thus, the MCI responds to both the number and severity of revealed preference violations. For example, if the MCI is 0.004, it means that the average cost of preference relations that must be removed to render the dataset consistent with GARP is 0.4 percent.
@@
To summarize, these indices measure the degree of GARP violations from different perspectives. More detailed theoretical discussions can be found in Apesteguia and Ballester (2015), Echenique (2021) and Polisson and Quah (2024).
3 Scanner Data and Experimental Design

### v2

@@ Abstract
Choice consistency with utility maximization, as a key assumption in economics, has been extensively used to evaluate decision quality of individuals and to predict real-world outcomes across different contexts. Here we investigate the generalizability of consistency measures derived from budgetary decisions in the lab-in-the-field experiment and purchasing decisions using supermarket scanner data. In the first study, we observe a lack of correlation between consistency scores derived from risky decisions in the experiment and those from supermarket food purchasing decisions. In the second study, we observe moderate correlations between experimental tasks and low to moderate correlations across purchasing categories and over time periods within the supermarket. Moreover, consistency in the two settings exhibits distinct predictive validity in predicting consumer behavior. These results suggest that choice consistency, as a measure of decision quality, may be better characterized as a multidimensional skill set rather than a single-dimensional ability.
Keywords: Consumer choice, utility maximization, revealed preference, lab-in-the-field experiment
JEL: C91, D81, D91
@@ 1 Introduction
Central to economic analysis is the assumption that a decision maker (DM) maximizes her utility function given her budget constraint. Revealed preference analysis characterizes the conditions under which a DM, based on a given choice dataset, indeed maximizes a well-behaved utility function (Samuelson, 1938; Afriat, 1967; Varian, 1982). It thus provides a powerful framework and serves as a textbook example of normative approaches on how decisions should be made. Going beyond normative approach, researchers have examined the descriptive validity of this framework, and used it to assess choice consistency in various settings (Crawford and De Rock, 2014; Chambers and Echenique, 2016). A growing body of literature has measured consistency scores in risky, intertemporal, and social budgetary decisions in experiments, as well as in purchasing decisions derived from expenditure surveys and from scanner data in supermarkets. [^1] For studies utilizing experimental data, see Ahn et al. (2014); Andreoni and Miller (2002); Choi et al. (2007); Echenique et al. (2023); Fisman et al. (2007); Halevy et al. (2018); Polisson et al. (2020); Dembo et al. (2024); Ellis and Freeman (2024), and for studies employing purchasing data, see Blundell et al. (2003, 2008); Crawford (2010); Cherchye et al. (2020); Dean and Martin (2016); Echenique et al. (2011). Consistency with utility maximization is often referred to as choice consistency or rationality. For simplicity, we use choice consistency throughout this paper.
Furthermore, it has been suggested that consistency scores reflect an individual’s capacity to make sound decisions. This notion has been proposed as a significant factor contributing to wealth disparities among individuals in the general population (Choi et al., 2014) and development gaps across countries (Cappelen et al., 2023). Choice consistency has also been used to assess the impact of financial constraints due to payday and loan (Carvalho et al., 2016, 2024), and to evaluate the effectiveness of policy interventions in education and poverty (Kim et al., 2018; Li et al., 2023).
@@
Building on these insights into choice consistency and its applicability across settings, we examine its empirical validity and test its generalizability—the extent to which DMs with high consistency scores in one setting also demonstrate high scores in another. More specifically, do DMs with high consistency scores as consumers in a supermarket environment also have higher scores as participants in a controlled experimental setting? Two related questions arise. Do DMs show correlated consistency scores across risk, social, and food-related decisions in the experiment? Do DMs exhibit correlated consistency scores across consumption categories and time periods in the supermarket? On the one hand, it is plausible that measures of choice consistency can be generalized across settings, as they may reflect a DM’s overall capacity of making good decisions. Conversely, it is also conceivable that these measures are setting-specific, as DM may possess different preference structures, face varying budget constraints, and have distinct experiences or decision rules across settings, leading to divergent scores inferred from choice data.
@@
We address these questions by combining experimental and scanner data. First, we examine consistency scores of the same individuals who make risky decisions in the lab-in-the-field experiment and make consumption decisions in the supermarket. Using the scanner data, we restrict our attention to 6,144 consumers of whom we have purchase records of meat and vegetables—the two most common consumption categories—over 24 consecutive months, so that we have sufficient power for the revealed preference analysis (see Section 3). We then successfully invite 1,055 of these 6,144 consumers to participate in a budgetary task (Experiment 1) in which participants make 22 portfolio decisions between two Arrow securities with different budget lines (Choi et al., 2007, 2014; Halevy et al., 2018; Kim et al., 2018). They need to allocate 100 experimental tokens between two accounts in which tokens are converted to cash with different exchange rates, and the amount in one of the two accounts will be paid out randomly with 50 percent probability. Afterwards, by extracting these 1,055 consumers from the scanner data and aggregating the consumption at the month level, we obtain the choice dataset in the supermarket (Scanner Dataset 1).
@@
We measure the choice consistency by evaluating the extent to which choices adhere to the Generalized Axiom of Revealed Preference (GARP), a necessary and sufficient condition whereby a dataset can be rationalized by a well-behaved utility function (in accordance with utility maximization). To assess how closely an individual’s choice dataset complies with GARP, we compute the commonly used critical cost efficiency index (CCEI, Afriat, 1972), as well as the Houtman-Maks index (HMI, Houtman and Maks, 1985); money pump index (MPI, Echenique et al., 2011); and minimum cost index (MCI, Dean and Martin, 2016). Based on Experiment 1 and Scanner Dataset 1, our primary finding is that consistency scores are uncorrelated between risky decisions in the experiment and food consumption decisions in the supermarket regardless of the index used. While this observation might appear to be unsurprising to some, it does contrast with our informal expert prediction. [^2] One of the authors conducted an informal expert prediction when presenting this paper at the Bounded Rationality in Choice Conference (BRIC, 2024, Venice) and the Decision: Theory, Experiments, and Applications Workshop (D-TEA, 2024, Paris). Among the approximately 50 attendees at each conference, majority predicted moderate to high correlations between the consistency scores of risky decisions and food decisions, while only 2 to 3 attendees anticipated the correlation to be close to zero.
@@
Second, we test whether the observed near-zero correlation is attributed to differences in the types of decision—risky decisions versus food consumption decisions—or due to the disparities between experimental and supermarket settings and conduct a second experiment (Experiment 2). We invite another 302 consumers to participate in three budgetary tasks, including the risk task used in Experiment 1, the social task (Andreoni and Miller, 2002; Fisman et al., 2007) and the food task (Chen et al., 2023; Harbaugh et al., 2001). We find that consistency scores are moderately correlated across these three tasks, where the magnitude of the correlations is the highest between risk and social tasks.
@@
Third, we also examine the correlation of consistency scores of choices within the supermarket using Scanner Dataset 2. The dataset consists of 822 consumers in different consumption categories, including meat, vegetables, fruits, and snacks, and 938 consumers in different time periods from 2018 to 2021. We find that the correlations of consistency scores are generally significantly positive between consumption categories and over different time periods within the supermarket, although the magnitude is relatively low.
@@
We conduct several robustness checks on the observed correlations. Specifically, we (1) utilize alternative indices for HMI, MPI, and MCI; (2) control for the power to detect GARP violations; (3) impose additional assumptions on preference structures through revealed preference analysis; and (4) assess the fit of preference structural estimations. Our findings remain consistent across these variations, with two notable exceptions. Namely, when we impose more assumptions on preference structures or evaluate the fit of structural estimations, the correlations within the scanner data significantly improve, both across categories and over time.
@@
Our results indicate that choice consistency exhibits multidimensional characteristics, with distinct factors affecting the degree of consistency measures across different decision environments, such as randomness in choice behaviors, budget constraints, formation of preferences, and the application of heuristic rules. In particular, we find that the choices made in supermarkets are more likely to come from random DMs compared to those made in experimental settings. Furthermore, the consistency of supermarket choices is influenced by various contextual factors, such as seasonality, time of day, and individual shopping experiences. In contrast, consistency in experimental choices is shaped by learning effects across rounds of decisions, the use of heuristic rules, and participants’ cognitive abilities. These distinct elements play crucial roles in how individuals navigate decision-making in the supermarket and in the experiment. To support this distinction, we further show that consistency in experimental data predicts the frequency and magnitude of discount usage in consumer’s purchasing choices, while consistency in supermarket data predicts the month-to-month volatility in shopping patterns. Taken together, we highlight the multifaceted nature of decision-making across settings.
@@
Our study adds to the literature on measuring consistency with utility maximization based on revealed preference analysis. In addition to theoretical and experimental studies, choice consistency measures have been proposed as proxies of decision-making quality and have been widely used in applied settings. These studies apply choice consistency measures to examine individual differences in wealth, occupation, and borrowing behavior of high-cost loans, policy interventions aimed at education and poverty reduction, as well as country-level variations in development (e.g., Banks et al., 2019; Cappelen et al., 2023; Carvalho et al., 2016, 2024; Choi et al., 2014; Kim et al., 2018; Li et al., 2023). In the literature, most studies focus on a specific type of decision in either the experiment or the field. [^3] An exception is Kim et al. (2018), in which they collect choice data on both risky and intertemporal decisions in the laboratory experiment to evaluate the effects of an education program. While their study does not examine the correlation between consistency scores of risky and intertemporal decisions, we use their data and compute the Spearman’s rank correlation to be 0.51, which is in line with the observed correlations across tasks in our Experiment 2. Our study is the first to examine the generalizability of consistency measures across settings, and contributes to a better understanding of the concept of utility maximization.
@@
Our study also contributes to the literature on the generalizability of choice behaviors across settings (Camerer, 2015; Levitt and List, 2007; Snowberg and Yariv, 2021; Chapman et al., 2024). A question that has attracted great attention is the generalizability of risk preferences. For example, Weber et al. (2002) show that risk attitudes vary between domains; these include financial, health/safety, recreational, ethical, and social decisions. Dohmen et al. (2011) find that different measures of risk attitudes are imperfectly correlated in different settings. Barseghyan et al. (2011) use three deductible choices made in the domain of auto and homeowner insurance to separately estimate the individual risk attitude using a structural approach, and reject the null of fully domain-general risk attitudes. Einav et al. (2012) provide evidence to support positive correlations of risk preferences within five employer-provider insurance coverage decisions, and weaker relationships between these insurance decisions and investment decisions with respect to a 401(k) plan. In addition to risk preferences, moderate or weak correlations between settings have also been reported in time preferences (Augenblick et al., 2015; Burks et al., 2012); social preferences (Bruhin et al., 2019; Fehr and Leibbrandt, 2011; Galizzi and Navarro-Martinez, 2019); and strategic sophistication (Georganas et al., 2015; Levitt et al., 2011; Rubinstein, 2016). These studies have made important contributions to the extent to which theory-guided experimental findings can be generalized and applied to real-world scenarios, which is essential for theoretical development, experimental design, and practical application.
@@
In this study, we examine choice consistency using risky, social, and food decisions in two experiments and consumption purchases in the supermarket. On the one hand, the lack of correlation between experimental and scanner data suggests that consistency scores revealed from choice behaviors in various settings may not solely reflect the general ability to make good decisions according to one’s preferences. On the other hand, we do observe moderate correlations across three tasks in the experiment and some weak to moderate correlations across categories and time in the supermarket, which indicate that consistency measures can be generalized within settings.
@@
In addition to measurement and identification issues, these findings support that choice consistency is multidimensional. This perspective resonates with developments in the literature on intelligence, which have evolved over decades from a one-dimensional understanding of intelligence (Spearman, 1904) to the recognition of two dimensions: fluid and crystallized intelligence (Cattell, 1943), and ultimately to the widely accepted view of multidimensional intelligence (McGrew, 2009; Horn and Cattell, 1966). In the economics literature, accumulative evidence has provided support for the notion that choice behaviors are interconnected and can be reduced to several underlying common factors (Chapman et al., 2023; Dean and Ortoleva, 2019; Stango and Zinman, 2023), and the ability to make good decisions is multidimensional (Deming, 2021; Cerigioni, 2021; Ilut and Valchev, 2023). Building on these studies, our study takes an initial step to address an important yet unexplored question on the generalizability of choice consistency, a central notion in economic analysis.
@@
The remainder of the paper is organized as follows. Section 2 presents the theoretical background of revealed preference analysis. Section 3 describes the scanner data and the experimental design. Section 4 presents our main results, and Section 5 examines the underlying mechanisms. Section 6 investigates the predictive validity of consistency scores across settings. Section 7 provides some concluding remarks.
@@ 2 Theoretical Background
[…]
@@ 2.1 Indices of Choice Consistency
Critical Cost Efficiency Index (CCEI) A popular approach to measure the departure from the maximization of utility is the critical cost efficiency index (CCEI) proposed by Afriat (1972). A DM has a CCEI $e\in\left[0,1\right]$ if $e$ is the largest number with a well-behaved $U$ that rationalizes the dataset for every $x^{t}\in\mathcal{X}:$
@@
A CCEI of 1 indicates passing GARP perfectly. A CCEI less than 1—say 0.95—indicates that there is a utility function for which the chosen bundle $x^{t}$ is preferred to any bundle that is more than 5 percent cheaper than $x^{t}$. Put differently, the CCEI can be viewed as the amount by which a budget constraint must be relaxed to remove all GARP violations, because the DM is spending more money than is required to achieve the utility targets (Polisson and Quah, 2024). Alternatively, it may be related to the behavioral notion of just-noticeable difference, the innate inability of DM to distinguish similar bundles (Dziewulski, 2020).
@@
Houtman-Maks Index (HMI) Houtman and Maks (1985) propose an alternative approach by measuring the maximum number of observations in the observed sample that are consistent with rational choices. For example, an HMI score of 0.941 indicates that the highest proportion of subset of choices consistent with GARP is 94.1 percent.
@@
Money Pump Index (MPI) Echenique et al. (2011) provide an index to measure the amount of money one can extract from an individual for each violation of GARP. Their index is based on the idea that an individual with a GARP violation can be exploited by an “arbitrager” as a “money pump”. The “arbitrager” can choose the allocation $x^{1}$ at price $p^{2}$ and the allocation $x^{2}$ at price $p^{1}$, then trade $x^{1}$ with the individual at $p^{1}$ and $x^{2}$ at $p^{2}$, which yields a profit of $p^{1}(x^{1}-x^{2})+p^{2}(x^{2}-x^{1})$. Given multiple violations of GARP, a money pump cost will be associated with each violation. Following Echenique et al. (2011), we use the mean money pump costs for cyclic sequences of allocations with the length of two. For example, an MPI score equal to 0.059 means on average 5.9 percent of expenditure can be exploited by an “arbitrager” from GARP violation.
[…]
@@
In summary, these indices measure the degree of GARP violations from different perspectives. More detailed theoretical discussions can be found in Apesteguia and Ballester (2015), Echenique (2021) and Polisson and Quah (2024).
3 Experimental Design and Scanner Data